[Hallc_running] [New Comment] RC Daily Update

rmontgom at jlab.org rmontgom at jlab.org
Fri Nov 23 16:15:01 EST 2018

 - by rmontgom on Fri, 11/23/2018 - 16:10.

pre-scale check - I added a comment to notices in effect for shift crews:
*please check also that the pre-scaler gui is active or if it needs refreshed


Logentry Text:
- Day shift yesterday continued production taking on Group 11 (Q2 = 5.0, x = 0.5, pi+).
- Swing shift yesterday completed Group 11, moved to Group 12 (which included a polarity change). A planned rate test at 12-1 was postponed to 12-3 due to trip rates/tune beam. There was a <1hr down time for an injector issue (DTM entry 5506).
- Owl crew continued data taking on group 12. During 12-3 a line voltage glitch at 04:51dropped beam to all halls. This was followed by several line voltage glitches (DTM entry 5508). This resulted in an SHMS HB interlock failure. Resetting interlock did not solve the issue, upon expert advice/approval the SHMS HB, Q3, dipole power supplies were power cycled via controlled access which solved the issue. Another access was required to open the SHMS shutter which closed during power cycle. This was followed by a fifth pass RF separator LCW flow issue for 48mins (DTM entry 5509).
- Day shift continue to take data on group 12.
- Group 12 will be followed by SHMS electron singles (details at https://hallcweb.jlab.org/wiki/index.php/CSV_Fall_2018_Run_Plan#Kinematic_Group_12:_Q2_.3D_5_GeV2.2C_x_.3D_0.5), then Group 8 (Q2 = 3.979, x = 0.5, pi-), then Group 7 (Q2 = 3.979, x = 0.5, pi+).

RC Meeting:
- We have no planned RC meetings until Monday. Unless there is a special request.

Notices in effect for shift crews:
* NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''
* Shift leaders, please fill out the check list once per shift!
* It is no longer required to fill out the standard.kinematics file online, it is now done automatically
* Split the large desired charge between multiple runs (~1 hour per run)
* Listed currents are estimates. For runs at less than 70uA, increase current until SHMS singles rates are between 450k-500k.
* Remember to check trigger pre-scales before runs (PS6=0 all the time, adjusting PS4 to give at least 100Hz, and no more than 200Hz, of HMS singles, PS1=PS2=PS3=PS5=-1). Check also that the pre-scaler gui is active and doesn't need refreshed.
* How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Monday, Nov 26: Energy change to 1.86 GeV/pass and move to 500 MHz laser. Hall will go to restricted access Monday morning and aim is to be locked up again Tuesday Swing. There has been a separate e-mail about shift updates. TO needed for all upcoming shifts. SL cancelled Mon day 11/26, Mon swing 11/26, Tue owl 11/27, Tue day 11/27. SL on standby Tue swing 11/27, Wed owl 11/28 and Wed day 11/28. SL shifts normal from Wed swing 11/28 currently. If on standby please be ready to come to counting house. We will know more Tue/Wed about expected beam delivery so stay tuned.

Planned accesses to the hall:
Monday, Nov 26: Work in the hall should be coordinated with Walter K. Important aerogel change taking place by V. Berdnikov (~8 hours duration Mon). HGC work will take place by B. Sawatzky and University of Regina team. D. Meekins will investigate the target loop 3 JT valve issue. V. Kumar (Regina group) walkthrough Monday morning (exact time TBC).
Tuesday, Nov 27: installation of brushes near the spectrometer wheels for S. Lassiter (originally planned for Tue/Wed).
Please let me know if there are any other plans for work in the hall Mon/Tue.

Special requests:
Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463). Currently we expect to delay this, to finish run plan before Mon.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3630349
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181123/d98b1559/attachment.html>

More information about the Hallc_running mailing list